OpenStudy (anonymous):

-(2x-6) expand brackets

OpenStudy (anonymous):

-1(2x-6)

OpenStudy (anonymous):

Using Distributive property you can do it.. Multiplication by minus sign here means multiplication by -1./.. so according to distributive property a*(x+y)=a*x+a*y also -1*-1 = 1 -1 * 1 = -1 1*1 = 1 ie -ve multiplied by -ve = +ve +ve multiplied by -ve = -ve -ve multiplied by +ve = -ve +ve multiplied by +ve = +ve hope now you can do it now by yourself

OpenStudy (anonymous):

Here * means Multiplication

OpenStudy (anonymous):

that is a*b = c means a is multiplied by b and the answer is c

OpenStudy (anonymous):

on multiplying -ve with -ve, u get +ve. -ve*+ve gives _ve.

OpenStudy (anonymous):

@Raeee_009 are you following?

OpenStudy (anonymous):

dont understand any of that

OpenStudy (anonymous):

ok..Let us start from the beginning... expand bracket means multiplying the terms outside the bracket with the terms inside.. that is for example 1.expand bracket a*(x+y) means multiply 'a' with the terms inside the bracket You understand that??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

well good.. so what are the terms inside the bracket in a*(x+y) ?

OpenStudy (anonymous):

6-2X

OpenStudy (anonymous):

axy?

OpenStudy (anonymous):

@Raeee_009 is it .. axy?? Here if we want to find the term inside bracket of blow question s*(t) inside the bracket term means the term inside the symbols ( ) here the term is side the symbol is 't' the term outside the bracket means the term outside the symbol ( ) here the term outside that symbol is s... understood?

OpenStudy (anonymous):

so if that is the case can you tell me so what is the term inside the bracket in a*(y) ?

hartnn (hartnn):

let me try with a diagram :